OpenStudy (anonymous):

For what values of p is y = x^p a solution to the differential equation x^2y'' - 3xy' - 5y = 0 ?

OpenStudy (anonymous):

Anyone?

myininaya (myininaya):

y=x^p y'=px^{p-1} y''=p(p-1)x^{p-2} x^2(p(p-1)x^{p-2})-3x(px^{p-1})-5(x^p) =p(p-1)x^p-3x^p -5x^p =(p^2-p)x^p-8x^p =(x^p)(p^2-p-8)

myininaya (myininaya):

do p^2-p-8=0 to find your p's

OpenStudy (anonymous):

let y=x^r y'=rx^r-1 y''=r(r-1)x^r-2 x^r(r-1)-3x^r(r)-5x^r=0 x^r(r^2-r-3r-5)=0 x^r(r^2-4r-5)=0 x^r((r-5)(r+1))=0 (r-5)(r+1)=0 therefore the real roots are y(x)= c1x^5+c2x^-1

myininaya (myininaya):

oops i missed something

OpenStudy (anonymous):

did you check mine myin

myininaya (myininaya):

nut got it

myininaya (myininaya):

i forgot my p in the 3x^p term

myininaya (myininaya):

suppose to be 3px^p

myininaya (myininaya):

gj nut

OpenStudy (anonymous):

ok ok thanks myin ;)

OpenStudy (anonymous):

Thanks a lot bnut.

OpenStudy (anonymous):

And thank you too myininaya.

myininaya (myininaya):

np